Rutro Corp. makes 59,000 units per year of a part it uses in the products it manufactures. The unit product cost of this part is computed as follows: Direct material $21.00 Direct labor 23.00 Variable manufacturing overhead 8.00 Fixed manufacturing overhead 30.00 Unit product cost $82.00 An outside supplier has offered to sell the company all of the 59,000 parts it needs for $75.00 a unit. If the company accepts this offer, the facilities now being used to make the part could be used to make more units of another product that is in high demand. The additional contribution margin on this other product would be $310,000 per year. If the part were purchased from the outside supplier, all of the direct labor costs of the part would be avoided. However, 64% of the fixed manufacturing overhead cost being applied to the part would continue even if the part were purchased from the outside supplier. This fixed manufacturing overhead cost would be applied to the company's remaining products. Required: How much of the unit product cost of $82.00 is relevant in the decision of whether to make or buy the part?

What is the net total dollar advantage (disadvantage) of purchasing the part rather than making it?

Should Rutro continue to manufacture the part or buy it?

What effect do fixed costs have in determining whether to continue manufacturing the part or buy it?

What would have to happen to fixed costs for your answer to change on the issue of continuing to manufacture the part or buy it?

What is the maximum amount the company should be willing to pay an outside supplier per unit for the part if the supplier commits to supplying all 59,000 units required each year?
Answer #1

Relevant cost per unit = avoidable cost

= 21+23+8+30*36%

=$62.8 per unit

advantage of purchasing = (Relevant cost of making- cost of buying)*number of units + additional contribution margin earned

=(62.8-75)*59000+310,000

=-$409,800 disadvantage

3. Continue to manufacture

4. Only avoidable fixed costs are relevant for the purpose of this decision as unavoidable fixed costs will be incurred in both the cases.

5. Fixed costs will become avoidable on buying

6. Maximum amount = avoidable cost + contribution margin earned per unit

= 62.8 + 310,000/59000

=$68.05 per unit